A hose ejects water at a speed of 20 cms-1 through a hole of area 100 cm2. If the water strikes a wall normally, calculate the force on the wall in newton, assuming the velocity of the water normal to the wall is zero after collision.

A  hose  ejects  water  at  a  speed  of  20  cms-1  through  a  hole  of  area  100  cm2.  If  the  water  strikes  a  wall normally,  calculate  the  force  on  the  wall in  newton,  assuming  the  velocity  of  the  water  normal  to  the wall  is  zero  after  collision. 

Grade:10

2 Answers

Ananya
2 years ago
Volume of water flowing out of the pipe per sec= 20*100=2000cm³
Density of water=1g/cm³
Mass of water flowing per sec= 2000g=2kg
F=ma
F=2(0-20)/1=-40N
Vibekjyoti Sahoo
2 years ago
Volume of water flowing out of the pipe per sec= 20*100=2000cm³
Density of water=1g/cm³
Mass of water flowing per sec= 2000g=2kg
F=ma
F=2(0-20)/1=-40N